Suppose capital share of output α increased suddenly and permanently in a small open
market economy. What happens to real wage rate, real rental rate, price level, real interest rate, real exchange
rate, consumption, and savings in long run and very-long run? Assume a Cobb Douglass production function.
1
Expert's answer
2020-12-24T13:17:17-0500
"U=L^ {\alpha}K^{1-\alpha}"

"\frac {\delta U}{\delta L}=\alpha (\frac{L}{K})^{\alpha-1}"


"\frac {\delta U}{\delta K}=(1-\alpha)(\frac {L}{K})^{\alpha}"